# Social Optimization in Noncooperative Games under Central Regulation

### v1

@@ 4           Update
See Appendix C. ∎
Corollary 3.
$\frac{1}{K}\sum_{k=0}^{K-1}\mathrm{E}[\|\nabla\hat{\mathbf{F}}(\theta_{k})\|^{2}]=\mathcal{O}(\frac{1}{\sqrt{K}}).$
Remark 4.
From Corollary 1, it is evident that in the inexact regime, with an appropriately selected inner iteration number $t_{k}$, our proposed algorithm can attain a sublinear convergence rate of $\mathcal{O}(\frac{1}{\sqrt{K}}+\frac{1}{K^{s}})$ with $s\in[0,1)$. On the other hand, one can see from Lemma 5 that an increasing number of iteration $t_{k}$ derives a more accurate inner loop solution, which ultimately results in a faster convergence speed owing to a larger $s$. Specially, the order of magnitude of (33) stands at $\mathcal{O}(\frac{1}{\sqrt{K}})$ when choosing $t_{k}\geq\lceil\frac{-s\ln(k+1)}{\ln q}\rceil$ and $s\in[\frac{1}{2},1)$. Moreover, according to Corollary 3, when the low-level game can be solved exactly, i.e., $\varepsilon_{k}=0$$\forall k=0,\ldots,K-1$, the convergence rate is also $\mathcal{O}(\frac{1}{\sqrt{K}})$. Note that the bound $\mathcal{O}(\frac{1}{\sqrt{K}})$ recovers the existing convergence rate of nonconvex nonsmooth optimization in [34]. Additionally, it should be emphasized that by utilizing the continuity of $x(\theta)$, if $\theta_{k}$ converges to the optimal solution $\theta^{*}\in\hat{\Theta}^{*}$, then $x(\theta)$ will correspondingly converges to $x(\theta^{*})$.
IV NUMERICAL SIMULATION
Consider the problem of flexible electric vehicle charging control [35]. Specifically, there are $N$ residents in a neighborhood who commute daily using electric vehicles and charge their vehicles at a public charging station. The charging process is regulated by a central authority (such as a power company or government) through adjustable price policies $\theta$ within a specified range $\Theta=[\theta_{\min},\theta_{\max}]$. Due to battery capacity limitations, every resident $i$ has its charging consumption $x_{i}\in X_{i}=[0,x_{i}^{\max}]$ and aims to minimize its electricity bill $f_{i}(x_{i},x_{-i},\theta):=c_{i}(x_{i}-d_{i})^{2}+a_{i}x_{i}+\lambda_{i}(x_{i}-\sum_{i=1}^{N}x_{i}/N)^{2}+rx_{i}\theta$. Herein, $c_{i},d_{i},a_{i},r>0$ are constants, and $\lambda_{i}$ captures resident $i$’s sensitivity to deviation from average consumption. Moreover, $x_{i}\theta$ represents the charging cost imposed by the central regulator when resident $i$ selects a specific $x_{i}$. In order to maintain a balanced demand on the power grid demand and reduce total consumption of social resources, the central authority interacts with residents and is interested in miniming the overall cost of all residents.
Set $N=10$. For every $i\in[10]$, set $c_{i}=4+i$, $d_{i}=7+2i$, $a_{i}=10+i$, $\lambda_{i}=0.1$, and $r=1$. Moreover, let set constraints $X_{i}=[0,25]^{10}$ and $\Theta=[1,3]$. Choose stepsizes $\alpha=0.00001$ and $\gamma=0.01$, smoothing parameter $\xi=0.0001$, and inner iteration number $t_{k}=\lceil 5\ln(k+1)\rceil$. The evolution of residual $|\theta_{k}-\theta^{*}|$ versus iteration $k$ is plotted in Fig. 2, demonstrating that the regulator can find the optimal decision successfully. Furthermore, Fig. 3 gives rises to the trajectory of $|\nabla\hat{\mathbf{F}}(\theta_{k})|$, from which it can be seen that the proposed algorithm achieves a sublinear convergence rate.
Figure 2: The trajectory of $|\theta_{k}-\theta^{*}|$.
Figure 3: The trajectory of $|\nabla\hat{\mathbf{F}}(\theta_{k})|$.
V CONCLUSION
In this work, a new framework inspired by noncooperative games, i.e., social optimization, has been explored, in which each cost function depends not only on all players’ strategies, but also on an intervention decision of a central regulator. This problem possesses a bilevel structure, where the low-level players aim to seek the NE associated with the regulator’s decision in a noncooperative game, and the goal of the high-level regulator is to minimize the sum of all players’ objectives to achieve the social optimum. To address this constrained, nonconvex and nonsmooth problem, a gradient-free method was developed by virtue of the randomized smoothness and Moreau smoothness techniques. It was proved that the designed algorithm converges at a sublinear rate, irrespective of whether the underlying game is solved exactly or inexactly. Potential avenues for future research may be on stochastic objective functions and incorporate equality and inequality constraints.
Appendix A Proof of Lemma 1
Proof.
which immediately leads to
$\langle-G(x(\theta),\theta),x(\theta^{\prime})-x(\theta)\rangle\leq 0,\ \langle-G(x(\theta^{\prime}),\theta^{\prime}),x(\theta)-x(\theta^{\prime})\rangle\leq 0.$
From the above two inequalities, one has
and hence, by Assumption 2-i) and 2-ii), it can be derived that
$\mu\|x(\theta^{\prime})-x(\theta)\|^{2}$ $\leq\langle G(x(\theta^{\prime}),\theta)-G(x(\theta),\theta),x(\theta^{\prime})-x(\theta)\rangle$
Therefore, it has
Subsequently, for any $\theta,\theta^{\prime}\in\mathbb{R}^{n}$, by applying (6), Assumption 1-iv), and (34), one can obtain
$|f_{i}(x(\theta),\theta)-f_{i}(x(\theta^{\prime}),\theta^{\prime})|$
$\leq(L_{x}\frac{l_{\theta}}{\mu}+L_{\theta})\|\theta-\theta^{\prime}\|.$
Consequently, the assertion (12) holds for $L_{F}=L_{x}\frac{l_{\theta}}{\mu}+L_{\theta}$. ∎
Appendix B Proof of Lemma 6
Proof.
$=\nabla\hat{F}(\theta_{k})+\nabla\hat{\chi}_{\Theta}(\theta_{k})$
$=\nabla\hat{\mathbf{F}}(\theta_{k}).$
It follows from the definitions of ^$\hat{\nabla}_{\varepsilon_{k}}F_{i}(\theta_{k})$ and ^$\hat{\nabla}F_{i}(\theta_{k})$, together with the triangle inequality and (6) that
^$\|\hat{\nabla}_{\varepsilon_{k}}F_{i}(\theta_{k})-\hat{\nabla}F_{i}(\theta_{k})\|$ $\leq\frac{n}{\xi}L_{x}\|x_{\varepsilon_{k}}(\theta_{k}+\xi u_{k})-x(\theta_{k}+\xi u_{k})\|$
$\quad+\frac{n}{\xi}L_{x}\|x_{\varepsilon_{k}}(\theta_{k})-x(\theta_{k})\|.$
Hence, it obtains that
^$\|\hat{\nabla}_{\varepsilon_{k}}F_{i}(\theta_{k})-\hat{\nabla}F_{i}(\theta_{k})\|^{2}\leq 4\frac{n^{2}}{\xi^{2}}L_{x}^{2}\varepsilon_{k}.$ (35)
Utilizing the $c_{r}$-inequality $\|\sum_{i=1}^{s}a_{i}\|^{2}\leq s\sum_{i=1}^{s}\|a_{i}\|^{2}$, it has
which yields (30) by combining with (35).
As for (31), note that
$=\mathrm{E}[\|\hat{\nabla}F(\theta_{k})\|^{2}|\mathscr{F}_{k}]+\|\nabla\hat{F}(\theta_{k})\|^{2}-2\mathrm{E}[\hat{\nabla}F(\theta_{k})^{\textit{T}}\nabla\hat{F}(\theta_{k})|\mathscr{F}_{k}]$
where the last equality holds due to
$\mathrm{E}[\hat{\nabla}F(\theta_{k})^{\textit{T}}\nabla\hat{F}(\theta_{k})|\mathscr{F}_{k}]$ $=\mathrm{E}[(\hat{\nabla}F(\theta_{k}))^{\textit{T}}|\mathscr{F}_{k}]\nabla\hat{F}(\theta_{k})$
$=\|\nabla\hat{F}(\theta_{k})\|^{2}.$
Recall from Lemma 1 that $F(\theta)$ is $NL_{F}$-Lipschitz continuous. Then (31) holds by (16) in Lemma 2. Therefore, the proof is completed. ∎
Appendix C Proof of Corollary 2
Proof.
@@ Appendix C Proof of Corollary 2
Applying the $\frac{nNL_{F}+1}{\xi}$-Lipschitz continuity of $\nabla\hat{\mathbf{F}}$ and substituting ^$\hat{\nabla}_{\varepsilon_{k}}F(\theta_{k})$ in (24) by ^$\hat{\nabla}F(\theta_{k})$, one can obtain that
$\hat{\mathbf{F}}(\theta_{k+1})$ $\leq\hat{\mathbf{F}}(\theta_{k})-\alpha\nabla\hat{\mathbf{F}}(\theta_{k})^{\textit{T}}(\hat{\nabla}F(\theta_{k})+\nabla\hat{\chi}_{\Theta}(\theta_{k}))$
@@
Taking conditional expectation with respect to $\mathscr{F}_{k}$ on both sides of (C) and relying on (29) in Lemma 6, one has
@@
The rest of the proof follows the same analysis as Theorem 1. ∎

### v2

@@ Appendix G Proof of Corollary 2
Applying the $\frac{nNL_{F}+1}{\xi}$-Lipschitz continuity of $\nabla\hat{\mathbf{F}}$ and substituting $\hat{\nabla}_{\varepsilon_{k}}F(\theta_{k})$ in (23) by $\hat{\nabla}F(\theta_{k})$, one can obtain that
$\hat{\mathbf{F}}(\theta_{k+1})$ $\leq\hat{\mathbf{F}}(\theta_{k})-\alpha\nabla\hat{\mathbf{F}}(\theta_{k})^{\top}(\hat{\nabla}F(\theta_{k})+\nabla\hat{\chi}_{\Theta}(\theta_{k}))$
@@
Taking conditional expectation with respect to $\mathscr{F}_{k}$ on both sides of (G) and relying on (28) in Lemma 6, one has
@@
Taking expectation on both sides of the preceding inequality, in view of (30) in Lemma 6 and invoking $\alpha\leq\frac{\xi}{2(nNL_{F}+1)}$, it derives that
$\mathrm{E}[\hat{\mathbf{F}}(\theta_{k+1})]$ $\leq\mathrm{E}[\hat{\mathbf{F}}(\theta_{k})]-\frac{1}{2}\alpha\mathrm{E}[\|\nabla\hat{\mathbf{F}}(\theta_{k})\|^{2}]$
$\quad+N^{2}n^{2}L_{F}^{2}\frac{nNL_{F}+1}{\xi}\alpha^{2}.$ (42)
For (G), summing over $k=0,\ldots,K-1$ and rearranging the inequality, it yields
$\frac{1}{K}\sum_{k=0}^{K-1}\mathrm{E}[\|\nabla\hat{\mathbf{F}}(\theta_{k})\|^{2}]$
$\leq\frac{2(\hat{\mathbf{F}}(\theta_{0})-\mathrm{E}[\hat{\mathbf{F}}(\theta_{K})])}{\alpha K}+2N^{2}n^{2}L_{F}^{2}\frac{nNL_{F}+1}{\xi}\alpha.$
The proof is completed by the definition of $\hat{\mathbf{F}}^{*}$. ∎